OpenStudy (anonymous):

Sharon is 18 and would like to buy a house when she is 34. What is the discount factor for today’s prices if the housing values increase 7% per year? 12.0% 14.29% 33.9% 66.1%

OpenStudy (anonymous):

I believe the answer is the third one. If housing values increased 7% per year for the next 34-18 = 16 years straight -- and, honestly, I am having a hard time not laughing hysterically as I type this, because that is the most absurd thing I've heard in a while -- then the price of a house will increase by (1.07)^16 = 2.952... That is, a house that costs $100,000 today will cost $295,216 in 16 years. Hence the discount factor is 1/(1.07)^16 = 0.3387... = 33.9%. That is, Sharon can buy a house today for 33.9% of what the same house will cost her when she is 34.